# A Library of LLM Intrinsics for Retrieval-Augmented Generation

### v1

@@ 3 Uncertainty Quantification
Granite 3.2 8b Instruct - Uncertainty Quantification is a LoRA adapter for ibm-granite/granite-3.2-8b-instruct, adding the capability to provide calibrated certainty scores when answering questions when prompted, in addition to retaining the full abilities of the ibm-granite/granite-3.2-8b-instruct model. The model is a LoRA adapter finetuned to provide certainty scores mimicking the output of a calibrator trained via the method in Shen et al. (2024).
@@ 3.1 Intended Use
Important note: Certainty is inherently an intrinsic property of a model and its abilities. Granite-3.2-8b-Uncertainty-Quantification is not intended to predict the certainty of responses generated by any other models besides itself or ibm-granite/granite-3.2-8b-instruct. Additionally, certainty scores are distributional quantities, and so will do well on realistic questions in aggregate, but in principle may have surprising scores on individual red-teamed examples.
@@ 3.1.1 Usage steps
See Section A.2 for an example describing how to use the Uncertainty Quantification intrinsic to answer questions and obtain intrinsic calibrated certainty scores.
@@ 3.1.2 Possible downstream use cases (not implemented)
  • RAG: Granite-3.2-8b-Uncertainty-Quantification is calibrated on document-based question answering datasets, hence it can be applied to giving certainty scores for answers created using RAG. This certainty will be a prediction of overall correctness based on both the documents given and the model’s own knowledge (e.g. if the model is correct but the answer is not in the documents, the certainty can still be high).
@@ 3.2 Evaluation
The model was evaluated on the MMLU [^3] https://huggingface.co/datasets/cais/mmlu datasets (not used in training). Shown are the Expected Calibration Error (ECE) [^4] https://towardsdatascience.com/expected-calibration-error-ece-a-step-by-step-visual-explanation-with-python-code-c3e9aa12937d for each task, for the base model (Granite-3.2-8b-instruct) and Granite-3.2-8b-Uncertainty-Quantification. The average ECE across tasks for our method is 0.064 (out of 1) and is consistently low across tasks (maximum task ECE 0.10), compared to the base model average ECE of 0.20 and maximum task ECE of 0.60. Note that our ECE of 0.064 is smaller than the gap between the quantized certainty outputs (10% quantization steps). Additionally, the zero-shot performance on the MMLU tasks does not degrade, averaging at 89%.
@@ 4 Hallucination Detection
Granite 3.2 8b Instruct - Hallucination Detection is a LoRA adapter for ibm-granite/granite-3.2-8b-instruct fine-tuned for the hallucination detection task of model outputs. Given a multi-turn conversation between a user and an AI assistant ending with an assistant response and a set of documents/passages on which the last assistant response is supposed to be based, the adapter outputs a hallucination risk range for each sentence in the assistant response.
@@ 4.1 Intended Use
  • conversation: A list of conversational turns between the user and the assistant, where each item in the list is a dictionary with fields role and content. The role equals to either user or assistant, denoting user and assistant turns, respectively, while the content field contains the corresponding user/assistant utterance. The conversation should end with an assistant turn and the text field of that turn should contain the assistant utterance with each sentence prefixed with a response sentence id of the form <rI>, where I is an integer. The numbering should start from 0 (for the first sentence) and be incremented by one for each subsequent sentence in the last assistant turn.
  • instruction: A task instruction, which is encoded as a dictionary with fields role and content, where role equals to system and content equals to the following string describing the hallucination detection task: "Split the last assistant response into individual sentences. For each sentence in the last assistant response, identify the faithfulness score range. Ensure that your output includes all response sentence IDs, and for each response sentence ID, provide the corresponding faithfulness score range. The output must be a json structure."
@@
Model output: When prompted with the above input, the model generates a range for faithfulness score (hallucination risk) for each sentence of the last assistant response in the form of a JSON dictionary. The dictionary is of the form {"<r0>": "value_0", "<r1>": "value_1", ...}, where each field <rI>, where I an integer, corresponds to the ID of a sentence in the last assistant response and its corresponding value is the range for faithfulness score (hallucination risk) of the sentence. The output values can show numeric ranges between 0-1 with increments of 0.1, where the higher values correponds to high faithfulness (low hallucination risk), and lower values corresponds to low faithfulness (high hallucination risk). Additionally, the model is trained to output unanswerable when the response sentence indicate that the question is not answerable, and to output NA when the faithfulness cannot be determined (ex: very short sentences).
See Section A.3 for an example describing how to use the Hallucination Detection intrinsic.
@@ 4.2 Evaluation
Hallucination Detection LoRA 67.6 77.4 72.2
@@ 4.3 Training Details
  • Faithfulness label generation: For creating the faithfulness labels for responses, we used the NLI-based technique available at Achintalwar et al. (2024).
@@
The LoRA adapter was fine-tuned using PEFT under the following regime: rank = 8, learning rate = 1e-5, and 90/10 split between training and validation.
5 Answerability Determination
Granite 3.2 8b Instruct - Answerability Determination is a LoRA adapter for ibm-granite/granite-3.2-8b-instruct fine-tuned for binary answerability classification task. The model takes as input a multi-turn conversation and a set of documents, and classifies whether the user’s final query is answerable or unanswerable based on the available information in the set of input documents.
5.1 Intended Use
See Section A.4 for an example describing how to use the Answerability Determination intrinsic.
5.2 Evaluation
5.2.1 Answerability Classification
    BigBird w/ MLP 49.2 68.5 57.3 48.0 29.2 36.3 46.8
    LLaMA 2-7B 72.2 71.0 71.6 71.4 72.6 72.0 71.8
    Granite 3.2-8b LoRA 84.2 68.0 75.2 73.1 87.2 79.5 77.4
BigBird w/ MLP 69.6 77.6 73.4 70.1 60.8 65.2 69.6
LLaMA 2-7B 86.9 89.4 88.2 87.3 84.5 85.9 87.1
Granite 3.2-8b LoRA 85.4 89.3 87.3 87.0 82.4 84.6 86.1
5.2.2 Comparing LoRA Adapter vs. Vanilla Granite for Answer Quality
We compare the performance of Granite 3.2-8b Instruct vs. Granite 3.2-8b LoRA adapter on a subset of MT-RAG Benchmark in Table 11. In this setup, each query is paired with only 5 retrieved passages as context. The true answerability label for each query indicates whether the query is answerable with respect to the retrieved context.
  • The RAGAS Faithfulness (RF) score (on truly answerable queries): This drops slightly with the LoRA adapter. However, this is not due to degraded generation quality, but rather because the model labels more truly answerable queries as unanswerable and abstains from answering.
  • Joint Answerability-Faithfulness Score (JAFS):
    The LoRA adapter achieves a 7% lift on this metric - rewarding the model for correctly abstaining on unanswerable queries and for being faithful when it chooses to answer.
Model Unans. F1 Ans. F1 Unans. Recall Ans. Recall RF (on Truly Answerable) JAFS
Granite 3.2-8b Instruct 14 76 8 97 75 50
Granite 3.2-8b LoRA 47 77 37 88 70 57
Table 11: Comparison of Granite 3.2-8B Instruct vs. LoRA Adapter on Answerability and Faithfulness metrics using MT-RAG Benchmark.
5.3 Training Details
@@ 6 Citation Generation
Granite 3.2 8b Instruct - Citation Generation is a RAG-specific LoRA adapter for ibm-granite/granite-3.2-8b-instruct fine-tuned for the citation generation task. Given a multi-turn conversation between a user and an AI assistant ending with an assistant response and a set of documents/passages on which the last assistant response is supposed to be based, the adapter generates citations for the last assistant response from the provided documents/passages. The LoRA adapter has the following features:
@@ 6.1 Intended Use
Model output: When prompted with the above input, the model generates the citations for each sentence of the last assistant response in the form of a JSON dictionary. The dictionary is of the form {"<r0>": ..., "<r1>": ..., ...}, where each field <rI>, with I an integer, corresponds to the ID of the corresponding sentence in the last assistant response and its value is a list of context sentence IDs corresponding to the sentence(s) in the input documents that support the particular response sentence.
See Section A.5 for an example describing how to use the Citation Generation intrinsic to generate citations for a given assistant response.
@@ 6.2.2 Evaluation on LongBench-Cite
Model generating response Model generating citations Longbench-Chat (en) MultifieldQA (en) HotpotQA GovReport
R P F1 R P F1 R P F1 R P F1
Llama-3.1-70B-Instruct Llama-3.1-70B-Instruct 27.0 34.4 26.1 46.1 63.3 49.7 34.0 39.4 30.2 55.0 77.5 62.0
Llama-3.1-70B-Instruct Granite-3.2-8B LoRA citations 61.9 68.6 62.0 71.2 84.1 74.3 66.8 73.3 65.4 70.3 83.6 75.4
@@ 7.1 Query Rewrite plus Answerability Determination
Experimental Setup. To benchmark the above flows, we use MT-RAG conversations and Elser for retriever. QR and AD is performed with QR LoRA adapter and AD LoRA adapter with Granite 3.2-8B Instruct, respectively. We set the number of retrieved passages to $5$ across different retrieval strategies. For generation, we send the Granite 3.2-8B Instruct model the following information: the entire conversation, top-$5$ retrieved passages, and the model’s default RAG instruction prompt.
@@ 7.1.1 Evaluation: Answerability Classification
Table 14 shows the $F1$ scores on the task of answerability classification of the 4 RAG flows described above. Using the AD intrinsic significantly improves performance on unanswerable queries ($F1$ score increases from 14 to 47). The QR intrinsic also has an effect (though it’s much smaller): increasing the Recall@5 performance of the retriever makes some questions more likely to be answerable ($F1$ score increases from 76 to 82).
@@ 8 Conclusion
In this paper we introduce a library of LLM intrinsics for RAG. The intrinsics currently implemented are Query Rewrite, Uncertainty Quantification, Hallucination Detection, Answerability Determination, and Citation Generation. They are released as LoRA adapters for ibm-granite/granite-3.2-8b-instruct on HuggingFace, as well as through the recommended implementations in Granite IO, accompanied in both places with documentation and code. All the models are publicly released under an Apache 2.0 license for both research and commercial use. We describe the intended usage, training details, and evaluations for each intrinsic. We also introduce the notion of Composite Intrinsics, and describe one particular composition in detail, including in-depth evaluation of the created flow.

### v2

@@ 7 Uncertainty Quantification
Granite 3.3 8b Instruct - Uncertainty Quantification is a LoRA adapter for ibm-granite/granite-3.3-8b-instruct, adding the capability to provide calibrated certainty scores when answering questions when prompted, in addition to retaining the full abilities of the ibm-granite/granite-3.3-8b-instruct model. The model is a LoRA adapter finetuned to provide certainty scores mimicking the output of a calibrator trained via the method in Shen et al. (2024).
@@ 7.1 Intended Use
Important note: Certainty is inherently an intrinsic property of a model and its abilities. Granite-3.3-8b-Uncertainty-Quantification is not intended to predict the certainty of responses generated by any other models besides itself or ibm-granite/granite-3.3-8b-instruct. Additionally, certainty scores are distributional quantities, and so will do well on realistic questions in aggregate, but in principle may have surprising scores on individual red-teamed examples.
@@ 7.1.1 Usage steps
Note that the model was trained with 1/3 each of the data having (a) no system prompt, (b) the default generic Granite system prompt, and (c) the default Granite RAG prompt. As a result, the model is robust to these choices of system prompt.
@@
For an example describing how to use the intrinsic, refer to the documentation of the RAG agent library at https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib.
@@ 7.1.2 Possible downstream use cases (not implemented)
  • RAG: Granite-3.3-8b-Uncertainty-Quantification is calibrated on document-based question answering datasets, hence it can be applied to giving certainty scores for answers created using RAG. This certainty will be a prediction of overall correctness based on both the documents given and the model’s own knowledge (e.g. if the model is correct but the answer is not in the documents, the certainty can still be high).
@@ 7.2 Evaluation
The model was evaluated on the MMLU [^5] https://huggingface.co/datasets/cais/mmlu datasets (not used in training). Shown are the Expected Calibration Error (ECE) [^6] https://towardsdatascience.com/expected-calibration-error-ece-a-step-by-step-visual-explanation-with-python-code-c3e9aa12937d for each task, for the base model (Granite-3.3-8b-instruct) and Granite-3.3-8b-Uncertainty-Quantification. The average ECE across tasks for our method is 0.064 (out of 1) and is consistently low across tasks (maximum task ECE 0.10), compared to the base model average ECE of 0.20 and maximum task ECE of 0.60. Note that our ECE of 0.064 is smaller than the gap between the quantized certainty outputs (10% quantization steps). Additionally, the zero-shot performance on the MMLU tasks does not degrade, averaging at 89%.
@@
System prompt and application point
Table 16 shows results of predicting the target certainty score before (Scenario 2) and after (Scenario 1) the assistant response under different prompting conditions. Note that the performance is robust to choice of system prompt (as all 3 are included in training). As should be expected, the Pre-answer (Scenario 2) performance is degraded, since it is predicting the certainty of the assistant answer before seeing the answer. That said, the MAE is still less than 1, i.e. less than the gap between adjacent certainty scores (integers 0 to 9).
Prompt Condition MAE Post Answer MAE Pre Answer
No system prompt 0.375 0.915
Standard system prompt 0.360 0.840
RAG prompt 0.385 0.840
Table 16: Mean Absolute Error (MAE) of predicting the target certainty score before (Scenario 2) and after (Scenario 1) the assistant response under different prompting conditions.
@@ 8 Hallucination Detection
Granite 3.3 8b Instruct - Hallucination Detection is a LoRA adapter for ibm-granite/granite-3.3-8b-instruct fine-tuned for the hallucination detection task of model outputs. Given a multi-turn conversation between a user and an AI assistant ending with an assistant response and a set of documents/passages on which the last assistant response is supposed to be based, the adapter outputs a hallucination risk for each sentence in the assistant response.
@@ 8.1 Intended Use
  • conversation: A list of conversational turns between the user and the assistant, where each item in the list is a dictionary with fields role and content. The role equals to either user or assistant, denoting user and assistant turns, respectively, while the content field contains the corresponding user/assistant utterance. The conversation should end with an assistant turn and the text field of that turn should contain the assistant utterance with each sentence prefixed with a response sentence id of the form <iI>, where I is an integer. The numbering should start from 0 (for the first sentence) and be incremented by one for each subsequent sentence in the last assistant turn.
  • instruction: A task instruction, which is encoded as a dictionary with fields role and content, where role equals to system and content equals to the following string describing the hallucination detection task: "Split the last assistant response into individual sentences. For each sentence in the last assistant response, identify the faithfulness by comparing with the provided documents and generate the faithfulness reasoning and faithfulness decision. Ensure that your output includes all response sentence IDs, and for each response sentence ID, provide the corresponding faithfulness reasoning and faithfulness decision. The output must be a json structure."
@@
Model output: When prompted with the above input, the model generates a json structure which contains a list of dictionaries, where the dictionaries in the list correspond to the sentences in the last assistant output. A dictionary contains 3 fields,
  1. 1.
    i - sentence id of the last assistant turn.
  2. 2.
    f - faithfulness/hallucination label for the sentence. This can take 4 values:
    1. (a)
      faithful - the sentence is faithful to the provided documents.
    2. (b)
      unfaithful - the sentence is not faithful to the provided documents.
    3. (c)
      partial - the sentence contain some claims that are not grounded to the contents of the documents.
    4. (d)
      NA - the sentence does not contain any claims.
  3. 3.
    r - reasoning for the the model’s faithfulness label.
For an example describing how to use the intrinsic, refer to the documentation of the RAG agent library at https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib.
@@ 8.2 Evaluation
Hallucination Detection LoRA 68.1 69.2 68.6
@@ 8.3 Training Details
  • Faithfulness label generation: For each turn of the multi-turn RAG conversations from the previous step, we used a multi-step synthetic faithfulness label and reasoning generation pipeline to generate faithfulness labels for the assistant response.
@@
The LoRA adapter was fine-tuned using PEFT under the following regime: rank = 16, learning rate = 1e-5, and 90/10 split between training and validation.
@@ 9 Citation Generation
Granite 3.3 8b Instruct - Citation Generation is a RAG-specific LoRA adapter for ibm-granite/granite-3.3-8b-instruct fine-tuned for the citation generation task. Given a multi-turn conversation between a user and an AI assistant ending with an assistant response and a set of documents/passages on which the last assistant response is supposed to be based, the adapter generates citations for the last assistant response from the provided documents/passages. The LoRA adapter has the following features:
@@ 9.1 Intended Use
Model output: When prompted with the above input, the model generates the citations for each sentence of the last assistant response in the form of a JSON array. The array is of the form [{"r": 0, "c": [...]}, {"r": 1, "c": [...]}, ...}], where a JSON object of the form {"r": k, "c": [l, m]}, where k, l, m integers, denotes that the response sentence with ID <rk> is supported by the context sentences with IDs <cl> and <cm>.
For an example describing how to use the intrinsic, refer to the documentation of the RAG agent library at https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib.
@@ 9.2.2 Evaluation on LongBench-Cite
Model generating response Model generating citations Longbench-Chat (en) MultifieldQA (en) HotpotQA GovReport AVG
R P F1 R P F1 R P F1 R P F1 F1
Llama-3.1-70B-Instruct Llama-3.1-70B-Instruct 27.0 34.4 26.1 46.1 63.3 49.7 34.0 39.4 30.2 55.0 77.5 62.0 42.0
Llama-3.1-70B-Instruct Granite-3.3-8B LoRA citations 57.6 60.3 58.4 71.5 82.5 75.0 65.3 71.3 63.8 72.8 83.5 77.2 68.6
@@
  • In addition to the evaluation results per dataset, we also report the AVG F1, computed as the average of the four dataset-specific F1 scores.
@@ 10.1 Query Rewrite plus Answerability Determination
Experimental Setup. To benchmark the above flows, we use MT-RAG conversations and Elser for retriever. QR and AD is performed with QR LoRA adapter and AD LoRA adapter with Granite 3.3-8B Instruct, respectively. We set the number of retrieved passages to $5$ across different retrieval strategies. For generation, we send the Granite 3.3-8B Instruct model the following information: the entire conversation, top-$5$ retrieved passages, and the model’s default RAG instruction prompt.
@@ 10.1.1 Evaluation: Answerability Classification
Table 20 shows the $F1$ scores on the task of answerability classification of the 4 RAG flows described above. Using the AD intrinsic significantly improves performance on unanswerable queries ($F1$ score increases from 17 to 69). The QR intrinsic also has an effect (though it’s much smaller): increasing the Recall@5 performance of the retriever makes some questions more likely to be answerable ($F1$ score increases from 77 to 83).
@@ 11 Conclusion
In this paper we introduce a library of LLM intrinsics for RAG. Six intrinsics are currently implemented by training a LoRA adapter are Query Rewrite, Context Relevance, Answerability Determination, Uncertainty Quantification, Hallucination Detection, and Citation Generation. These have been released on HuggingFace as the Granite 3.3 8b RAG Agent Library, available at: https://huggingface.co/ibm-granite/granite-3.3-8b-rag-agent-lib. Two additional intrinsics, Query Expansion and Passage Reranking, are pure software-based implementations and are only available through Granite IO: https://github.com/ibm-granite/granite-io. All the models are publicly released under an Apache 2.0 license for both research and commercial use. We describe the intended usage, training details, and evaluations for each intrinsic. We also introduce the notion of Composite Intrinsics, and describe one particular composition in detail, including in-depth evaluation of the created flow.